# Gated InAs quantum dots embedded in surface acoustic wave cavities for low-noise optomechanics

### v1

@@
Gated InAs quantum dots embedded in surface acoustic wave cavities for low-noise optomechanics
@@
[…]
journal: opticajournalarticletype: Research Article{abstract*}
Self-assembled InAs quantum dots (QDs) are promising optomechanical elements due to their excellent photonic properties and sensitivity to local strain fields. Microwave-frequency modulation of photons scattered from these efficient quantum emitters has been recently demonstrated using surface acoustic wave (SAW) cavities. However, for optimal performance, a gate structure is required to deterministically control the charge state and reduce charge noise of the QDs. Here, we integrate gated QDs and SAW cavities using molecular beam epitaxy and nanofabrication. We demonstrate that with careful design of the substrate layer structure, integration of the two systems can be accomplished while retaining the optimal performance of each subsystem. These results mark a critical step toward efficient and low-noise optomechanical systems for microwave-to-optical quantum transduction.
Among various approaches for quantum transduction, hybrid systems involving mechanical resonators (confined phonons) are promising [10]. These electro-optomechanical systems take advantage of the slower speed of phonons compared to photons, resulting in phonons at gigahertz (GHz) frequencies with wavelengths comparable to optical photons. This is key for mediating interactions across the energy gap. It has been shown that phonons interact well with various types of quantum systems. For example, they piezoelectrically interact with superconducting qubits [11] and parametrically modulate various optical systems such as optical cavities [12] and single photon emitters (SPEs) in 2D materials [13, 14, 15], defect centers [16, 17, 18], and semiconductor quantum dots (QDs) [19, 20, 21, 22, 23]. However, integrating two standalone electromechanical and optomechanical systems is challenging due to the added system complexity and possible incompatibility [24].
In this paper, we address this integration challenge in an electro-optomechanical system involving gated InAs QDs and surface acoustic wave (SAW) resonator phonons. Our system combines peak optical performance in gated InAs QDs with state-of-the-art surface acoustic wave (SAW) resonators on GaAs. When combined with superconducting qubits[25, 26], these results pave the way for low-noise and efficient conversion of quantum information between microwave and optical domain.
Self-assembled InAs QDs have emerged as highly favorable SPEs and offer many benefits over other systems, such as trapped ions, color centers, vacancies, or defects. For example, previous work has demonstrated state-of-the-art end-to-end photon collection efficiencies of 57% [27], two-photon interference visibilities of about 99% [28], and repetition rates of approximately 1 GHz [27]. InAs QDs are embedded in semiconductor materials and are compatible with various fabricated photonic structures for efficient photon extraction [29, 30, 31, 32]. Their electrostatic environment and charge state can be controlled by simple gate structures [33]. With deterministic trapping of a single charge or hole, InAs QDs enable spin-photon entanglement [34] and remote spin-spin entanglement [35], which are crucial for quantum information processing applications. Finally, they exhibit a relatively simple energy level structure, approximating a two-level system, allowing for deterministic excitation with very simple optical pumping schemes. This provides a more promising route to generate single phonons that are important for phononic quantum technologies. In contrast, atoms and ions require active trapping mechanisms and complex optical pumping schemes for cooling and state preparation [36], and suffer from poor photon collection efficiencies into fiber [37]. Color centers, vacancies and defects are other solid-state SPEs with promising properties, but they suffer from non-unity quantum efficiencies and often fast phonon-assisted emission pathways [38].
[…]
@@
Figure 1: Illustration and micrographs of the integrated device. a) Optical micrograph of the fabricated device integrating gated QDs with SAWs. CPW: Co-planar waveguide. IDT: Interdigitated transducers. b) Illustration of the cross-sections A and B, corresponding to the horizontal and vertical white dashed lines in panel a. The main integration challenge is the loss introduced to the launching via the IDT ($\kappa_{IDT}$) and propagation of the SAW due to the doped GaAs layer below surface ($\kappa_{SAW}$). The microwave transmission through the coplanar waveguide will also be affected, introducing an effective loss $\kappa_{EM}$. As shown in b), the doped layer is etched away outside the black dashed region indicated in panel a) to reduce extra loss introduced to the microwave transmission line. A thin film of Ti/Au is evaporated inside the SAW resonator forming the top gate for the Schottky structure. An ohmic contact is made to the n-doped GaAs layer by Ni and AuGe, serving as the back gate. c) Scanning electron microscope (SEM) image of the Al IDT fingers. d) SEM image of the etched SAW DBR mirror. A slight curvature is added to reduce phonon diffraction loss.
@@
$\Delta V_{s}=\frac{\alpha^{2}/2}{1+\sigma^{2}_{xx}/\sigma^{2}_{m}},\quad\kappa=\frac{\alpha^{2}\sigma_{xx}/2\sigma_{m}}{1+\sigma^{2}_{xx}/\sigma^{2}_{m}}$ (1)
@@
Finally, the n-doped layer will also affect the microwave power delivered to the IDT via the co-planar waveguide (CPW), $\kappa_{EM}$. Due to the proximity of the n-doped layer to the surface (several hundred nanometers), the microwave mode in the CPW (with dimensions on the order of a few micrometers) is modified, thus reducing the characteristic transmission line impedance. The deviation of the impedance from a 50-ohm transmission line will reflect microwave signals before reaching the device, and further reduce the overall SAW launching efficiency. The solution for this is simple: the n-doped layer is unnecessary below the CPW structure and can be removed under the microwave circuits to recover the transmission line impedance. As shown in Fig 1b), the buried n-doped layer exists only in the region within the black dotted line, and is etched away everywhere else.
@@
Figure 2: Microwave characterization of the SAW delay lines and resonators. a) Reflection and b) transmission spectra of SAW delay lines with a 400 $\mu\text{m}$ gap. The similar peak transmission on substrates with buried n-doped GaAs compared to the same device on bulk GaAs shows the loss introduced by SAW launching and propagation is negligible with the proper choice of the depth and doping concentration of the n-doped GaAs layer. c) Peak transmitted SAW signal on the two substrates, varying the gap between the IDTs of the delay lines. The similar slope proves that the propagation loss introduced by the n-doped GaAs layer can be neglected. The error bars represent the noise in the transmission spectra as a result of the beating between the acoustic signal and electrical crosstalk. d) Reflection spectrum of a SAW resonator fabricated on a substrate with a buried n-doped GaAs layer, showing a Q factor of approximately 28,000.
@@
To experimentally verify these proposed solutions, we first fabricated acoustic delay lines formed by two IDTs separated by a gap of approximately 400 $\mu$m on a “charge-control substrate” (i.e., with the buried n-doped layer). We compare the delay lines’ performance to identical devices fabricated on mechanical-grade semi-insulating GaAs substrates without the n-doped layer. The microwave reflection and transmission of the acoustic delay lines are measured using a vector network analyzer (VNA) in a probe station at room temperature. As shown in Fig 2a), similarity in the reflection spectrum ($|S_{11}|^{2}$) dip sizes between identical devices on the two substrates indicates negligible reduction in piezoelectric coupling between IDTs and SAWs. SAW generation and attenuation are analyzed by looking at peak transmitted signals in the transmission spectrum ($|S_{21}|^{2}$) (Fig 2b) with different delay line gaps (Fig 2c). Variations in the peak transmitted power with propagation length directly reveal SAW propagation loss rates. The error bars in Fig 2c) represent the noise in the transmission spectra as a result of the beating between the acoustic signal and electrical crosstalk. Due to the buried n-doped layer, the electrical crosstalk is larger in the “charge control substrate”, leading to slightly higher noise level. However, similar transmission peak power and loss per unit length show that (1) microwave power is delivered equally well to the IDTs in both systems and (2) propagation loss introduced by the doped layer is not evident above our measurement uncertainty. To further verify the low loss introduced by the n-doped layer, we fabricated planar SAW resonators on charge-control substrates and characterized the SAW cavity quality factors Q with devices cooled to 1.7 K. By fitting the microwave reflection spectra, we extract intrinsic Q factors of approximately 28,000 (Fig. 2d). Remarkably, this is significantly better than Q factors derived from cavities with similar design parameters without conductive layers from previous publications [40]. This proves that the n-doped GaAs layer can be integrated with the SAW cavity without degrading electromechanical performance.
@@
Figure 3: Optical characterization of the hybrid device. a) PL spectrum of the gated QD, measured with spectrometer and CCD camera. Three charge plateaus are visible when varying bias voltage applied to the gate. b) Resonance fluorescence signal of the gated QD, measured with a superconducting nanowire single photon detector (SNSPD). A near-transform-limited linewidth is observed, which is a result of reduced charge noise by the gate structure. c) PL spectrum of the gated QD with SAW modulation at a fixed frequency (3.53388 GHz), measured with a Fabry-Perot filter and SNSPD. Both the QD’s resonant peak and sideband peaks are shifted by the Stark effect, with a rate of approximately 0.13 GHz/mV. d) PL spectrum of the gated QD at fixed gate bias with varying microwave drive frequency and constant microwave power, measured with a Fabry-Perot filter and SNSPD. A Lorentzian fit to the square of the modulation index ($\delta^{2}$) shows an acoustic resonance linewidth of approximately 232 kHz, corresponding to a Q factor approximately 15,000.
@@
After addressing each challenge individually, we proceed with fabrication of the proposed hybrid device. Fig 1a) shows an optical microscope image of the fully integrated device. The complete system contains a top gate of a thin (3nm/7nm) semi-transparent layer of Ti/Au inside the acoustic resonator, forming a Schottky diode structure. Optical measurements are performed on QDs directly below this thin metal gate. The thin metal layer is connected to a thick Ti/Au (10nm/300nm) layer outside the acoustic resonator for wirebonding (top right in Fig 1a). The metal surrounding the SAW resonator provides ohmic contact to the n-doped layer and ensures a uniform electric potential across the layer.
@@
To better assess the performance of the QDs, we perform resonance fluorescence measurements. These measurements are very sensitive probes of the QD’s linewidth and additional noise sources in the QD’s environment [49, 27]. We resonantly pump the QDs and reject reflected pump photons with a cross-polarization scheme, while fixing the bias voltage at one of the charge plateaus. The experimental setup is similar to that of Ref [20]. An exemplary resonance fluorescence spectrum is shown in Fig 3b). The measured optical linewidth is approximately 643.6 MHz, comparable to the state-of-the-art performance [27]. This narrow linewidth is a result of the reduced charge-noise of the gated QDs, compared to typical linewidth of 1.7 GHz without a gate structure.
@@
The gated QDs also interact with the SAW resonators optomechanically. We resonantly drive the SAW cavity with an external microwave source, and measure the PL of the QD as a function of bias voltage. The PL spectrum is measured with a tunable Fabry-Perot filter (600 MHz linewidth). As shown in Fig 3c), the QD is phase-modulated by the SAW field, resulting in a series of optical sidebands. The entire phase-modulated spectrum is Stark shifted by varying the bias voltage, with a rate approximately 0.13 GHz/mV, verifying that both QD charge control and SAW cavity can operate simultaneously.
@@
The modulated QD spectrum depends on the SAW field amplitude and can act as a probe to measure the acoustic Q factor of the fully integrated device under optical pumping [40]. We measure the QD PL spectrum with a fixed bias voltage and fixed microwave drive power while sweeping the microwave drive frequency around the SAW cavity mode. We then extract the modulation index, $\delta$, by fitting the phase-modulated PL spectrum as a function of the microwave drive frequency. A Lorentzian fit to $\delta^{2}$ shows a clear acoustic resonance with a linewidth of approximately 232 kHz, corresponding to a Q factor of approximately 15,000. Microwave refection measurement of the same cavity mode shows a Q factor of approximately 20,000. The discrepancy is likely due to heating of the cavity introduced by the pump laser, which can be avoided when pumping the QDs resonantly. This suboptimal Q factor of 20,000, compared to 28,000 shown in Fig 2d), is likely due to design imperfections in our cavities when introducing the thin Ti/Au layer. These design issues can be solved by more carefully compensating for SAW phase-velocity changes under the Ti/Au layer. Nonetheless, this Q factor is still comparable to those on substrates without an n-type GaAs layer [40]. Therefore, the integration of gated QDs and SAW resonators can be done without sacrificing the performance of individual components.

### v2

@@
Abstract
Self-assembled InAs quantum dots (QDs) are promising optomechanical elements due to their excellent photonic properties and sensitivity to local strain fields. Microwave-frequency modulation of photons scattered from these efficient quantum emitters has been recently demonstrated using surface acoustic wave (SAW) cavities. However, for optimal performance, a gate structure is required to deterministically control the charge state and reduce charge noise of the QDs. Here, we integrate gated QDs and SAW cavities using molecular beam epitaxy and nanofabrication. We demonstrate that with careful design of the substrate layer structure, integration of the two systems can be accomplished while retaining the optimal performance of each subsystem. These results mark a critical step toward efficient and low-noise optomechanical systems that truly leveraging the excellent properties of semiconductor QDs.
Introduction
Among various approaches for quantum transduction, hybrid systems involving mechanical resonators (confined phonons) are promising [10]. These electro-optomechanical systems take advantage of the slower speed of phonons compared to photons, resulting in phonons at gigahertz (GHz) frequencies with wavelengths comparable to optical photons. This is key for mediating interactions across the energy gap. It has been shown that phonons interact well with various types of quantum systems [11]. For example, they piezoelectrically interact with superconducting qubits [12] and parametrically modulate various optical systems such as optical cavities [13] and single photon emitters (SPEs) in 2D materials [14, 15, 16], defect centers [17, 18, 19], and semiconductor quantum dots (QDs) [20, 21, 22, 23, 24]. However, integrating two standalone electromechanical and optomechanical systems is challenging due to the added system complexity and possible incompatibility [25].
In this paper, we address this integration challenge in an electro-optomechanical system involving gated InAs QDs and SAW resonator phonons. Our system combines excellent optical performance in gated InAs QDs with state-of-the-art surface acoustic wave (SAW) resonators on GaAs, fully incorporating microwave and DC electronics, optics, and mechanics in a single platform. In general, this development is crucial to fully realize the advantages of QDs as optical two-level systems for optomechanics, including single-phonon generation [26] and acousto-optic preparation of exciton or biexciton states [27] which have so far been only theoretically proposed. In addition, interactions between acoustics and other types of SPEs (e.g., diamond vacancies) have recently caught attention as useful and interesting systems [28, 29, 30]. This work brings semiconductor QDs, with their many desirable properties, up to par with these other systems. When combined with superconducting qubits [31, 32], these results pave the way for low-noise and efficient conversion of quantum information between microwave and optical domain.
Self-assembled InAs QDs have emerged as highly favorable SPEs and offer many benefits over other systems, such as trapped ions, color centers, vacancies, or defects. For example, previous work has demonstrated state-of-the-art end-to-end photon collection efficiencies of 57% [33], two-photon interference visibilities of about 99% [34], and repetition rates of approximately 1 GHz [33]. InAs QDs are embedded in semiconductor materials and are compatible with various fabricated photonic structures for efficient photon extraction [35, 36, 37, 38]. Their electrostatic environment and charge state can be controlled by simple gate structures [39]. With deterministic trapping of a single charge or hole, InAs QDs enable spin-photon entanglement [40] and remote spin-spin entanglement [41], which are crucial for quantum information processing applications. Finally, they exhibit a relatively simple energy level structure, approximating a two-level system, allowing for deterministic excitation with very simple optical pumping schemes. This provides a promising route to generate single phonons that are important for phononic quantum technologies [26]. In contrast, atoms and ions require active trapping mechanisms and complex optical pumping schemes for cooling and state preparation [42], and suffer from poor photon collection efficiencies into fiber [43]. Color centers, vacancies and defects are other solid-state SPEs with promising properties, but they suffer from non-unity quantum efficiencies and often fast phonon-assisted emission pathways [44].
@@ Introduction
Figure 1: Illustration and micrographs of the integrated device. a) Optical micrograph of the fabricated device integrating gated QDs with SAWs. CPW: Co-planar waveguide. IDT: Interdigitated transducers. b) Illustration of the cross-sections A and B of the proposed design, corresponding to the horizontal and vertical red dashed lines in panel a. The main integration challenge is the loss introduced to the launching via the IDT ($\kappa_{IDT}$) and propagation of the SAW due to the doped GaAs layer below surface ($\kappa_{SAW}$). The microwave transmission through the coplanar waveguide will also be affected, introducing an effective loss $\kappa_{EM}$. As shown in b), the doped layer is etched away outside the black dashed region indicated in panel a) to reduce extra loss introduced to the microwave transmission line. A thin film of Ti/Au is evaporated inside the SAW resonator forming the top gate for the Schottky structure. An ohmic contact is made to the n-doped GaAs layer by Ni and AuGe, serving as the back gate. c) Scanning electron microscope (SEM) image of the Al IDT fingers. d) SEM image of the etched SAW DBR mirror. A slight curvature is added to reduce phonon diffraction loss.
Results
Integration Challenges
@@ Results
$\Delta V_{s}=\frac{\alpha^{2}/2}{1+\sigma^{2}_{xx}/\sigma^{2}_{m}},\quad\kappa=\frac{\alpha^{2}q\sigma_{xx}/2\sigma_{m}}{1+\sigma^{2}_{xx}/\sigma^{2}_{m}}$ (1)
@@
Finally, the n-doped layer will also affect the microwave power delivered to the IDT via the co-planar waveguide (CPW), $\kappa_{EM}$. Due to the proximity of the n-doped layer to the surface (several hundred nanometers), the microwave mode in the CPW (with dimensions on the order of a few micrometers) is modified, thus reducing the characteristic transmission line impedance from approximately 50 $\Omega$ to 3 $\Omega$. The deviation of the impedance will reflect approximately 88% of the microwave signals before reaching the device, and further reduce the overall SAW launching efficiency. The solution for this is simple: the n-doped layer is unnecessary below the CPW structure and can be removed under the microwave circuits to recover the transmission line impedance. As shown in Fig 1b), the buried n-doped layer exists only in the region within the black dotted line, and is etched away everywhere else.
@@
Figure 2: Microwave characterization of the SAW delay lines and resonators. a) Reflection and b) transmission spectra of SAW delay lines with a 400 $\mu\text{m}$ gap. The similar peak transmission on substrates with buried n-doped GaAs compared to the same device on bulk GaAs shows the loss introduced by SAW launching and propagation is negligible with the proper choice of the depth and doping concentration of the n-doped GaAs layer. c) Peak transmitted SAW signal on the two substrates, varying the gap between the IDTs of the delay lines. The similar slope proves that the propagation loss introduced by the n-doped GaAs layer can be neglected. The error bars represent the noise in the transmission spectra as a result of the beating between the acoustic signal and electrical crosstalk. d) Reflection spectra of SAW resonators with the same design parameters, fabricated on substrates with (bottom panel) and without (top panel) the buried doped layer, showing similar internal quality factors $Q_{i}$$\approx$ 28,000. Circle: data. Solid line: fit according to the model described in the Supplementary Information.
Microwave Measurements Characterizing the SAW Resonator
@@
To experimentally verify these proposed solutions, we first fabricated acoustic delay lines formed by two IDTs separated by a gap of approximately 400 $\mu$m on a “charge-control substrate” (i.e., with the buried n-doped layer). We compare the delay lines’ performance to identical devices fabricated on mechanical-grade semi-insulating GaAs substrates without the n-doped layer. The microwave reflection and transmission of the acoustic delay lines are measured using a vector network analyzer (VNA) in a probe station at room temperature. As shown in Fig 2a), similarity in the reflection spectrum ($|S_{11}|^{2}$) dip sizes between identical devices on the two substrates indicates negligible reduction in piezoelectric coupling between IDTs and SAWs. SAW generation and attenuation are analyzed by looking at peak transmitted signals in the transmission spectrum ($|S_{21}|^{2}$) (Fig 2b) with different delay line gaps (Fig 2c). Variations in the peak transmitted power with propagation length directly reveal SAW propagation loss rates. The error bars in Fig 2c) represent the noise in the transmission spectra as a result of the beating between the acoustic signal and electrical crosstalk. Due to the buried n-doped layer, the electrical crosstalk is larger in the “charge control substrate”, leading to slightly higher noise level. However, similar transmission peak power and loss per unit length show that (1) microwave power is delivered equally well to the IDTs in both systems and (2) propagation loss introduced by the doped layer is not evident above our measurement uncertainty. To further verify this, we fabricated and characterized losses and external coupling rates in planar SAW resonators on charge-control substrates with devices cooled to 1.7 K. By fitting the microwave reflection spectra (Supplementary Information), we extract an internal (external) quality factor $Q_{i}$ ($Q_{e}$) of approximately 28,000 (112,500) (Fig. 2d, lower panel). Both internal losses ($Q_{i}$) and electromechanical coupling ($Q_{e}$) are comparable to similar devices fabricated on a semi-insulating GaAs substrate (Fig. 2d, upper panel). In addition, the internal quality factor $Q_{i}$ is comparable with the state-of-the-art SAW cavities fabricated on a variety of materials at GHz frequencies [55, 56, 57, 58, 59, 60]. This proves that the n-doped GaAs layer can be integrated with the SAW cavity without degrading electromechanical performance. By comparing these results with previous results in the literature [46], it becomes clear that losses are dominated by features unrelated to the conductive layer, but rather to bulk scattering from the mirrors and IDTs. In particular, increased $Q_{i}$s and better coupling here, compared to Ref. [46], are related to the lighter IDT metal (Al compared to Nb or Au; Supplementary Information).
@@
Figure 3: Optical characterization of the hybrid device. a) PL spectrum of the gated QD, measured with spectrometer and CCD camera. Two charge plateaus are visible when varying bias voltage applied to the gate. b) Resonance fluorescence signal of the gated QD, measured with a superconducting nanowire single photon detector (SNSPD). A near-transform-limited linewidth is observed, which is a result of reduced charge noise by the gate structure. c) PL spectrum of the gated QD with SAW modulation at a fixed frequency (3.53388 GHz), measured with a Fabry-Perot filter and SNSPD. Both the QD’s resonant peak and sideband peaks are shifted by the Stark effect, with a rate of approximately 0.13 GHz/mV. d) PL spectrum of the gated QD at fixed gate bias with varying microwave drive frequency and constant microwave power, measured with a Fabry-Perot filter and SNSPD. A Lorentzian fit to the square of the modulation index ($\delta^{2}$) shows an acoustic resonance linewidth of approximately 232 kHz, corresponding to a Q factor approximately 15,000.
Optical Characterization of the Hybrid Device
@@
After addressing each challenge individually, we proceed with fabrication of the proposed hybrid device. Fig 1a) shows an optical microscope image of the fully integrated device. The complete system contains a top gate of a thin semi-transparent layer of Ti/Au inside the acoustic resonator, forming a Schottky diode structure. Optical measurements are performed on QDs directly below this thin metal gate. The thin metal layer is connected to a thick Ti/Au layer outside the acoustic resonator for wirebonding (top right in Fig 1a). The metal surrounding the SAW resonator provides ohmic contact to the n-doped layer and ensures a uniform electric potential across the layer.
@@
Resonance fluorescence is a very sensitive probe of the QD’s linewidth and additional noise sources in the QD’s environment [61, 33]. We resonantly pump the QDs and reject reflected pump photons with a cross-polarization scheme, while fixing the bias voltage at one of the charge plateaus. The experimental setup is similar to that of Ref [21]. An exemplary resonance fluorescence spectrum of the unmodulated QD is shown in Fig 3b). The measured optical linewidth is approximately 643.6 MHz, comparable to the state-of-the-art performance [33]. This narrow linewidth is a result of the reduced charge-noise of the gated QDs, compared to typical linewidth of approximately 1.7 GHz without a gate structure.
@@
The gated QDs also interact with the SAW resonators optomechanically. We resonantly drive the SAW cavity with an external microwave source, and measure the PL of the QD as a function of bias voltage. The PL spectrum is measured with a tunable Fabry-Perot filter (600 MHz linewidth). As shown in Fig 3c), the QD is phase-modulated by the SAW field, resulting in a series of optical sidebands. The entire phase-modulated spectrum is Stark shifted by varying the bias voltage, with a rate approximately 0.13 GHz/mV, verifying that both QD charge control and SAW cavity operate simultaneously.
@@
The modulated QD spectrum depends on the SAW field amplitude and can act as a probe to measure the acoustic $Q$ factor of the fully integrated device under optical pumping [46]. The acoustic Q factor sampled by the QD includes both internal and external loss channels. We measure the QD PL spectrum with a fixed bias voltage and fixed microwave drive power while sweeping the microwave drive frequency around the SAW cavity mode. We then extract the modulation index, $\delta$, by fitting the phase-modulated PL spectrum as a function of the microwave drive frequency. A Lorentzian fit to $\delta^{2}$ shows a clear acoustic resonance with a linewidth of approximately 232 kHz, corresponding to a Q factor of approximately 15,000. Microwave reflection measurement of the same cavity mode shows a total $Q$ of approximately 17,800 ($Q_{i}\approx$ 18,500; Supplementary Information). This small discrepancy may be due to local heating or additional charges in the cavity introduced by the pump laser, which can be avoided when pumping the QDs resonantly. This suboptimal internal quality factor of 18,500, compared to 28,000 shown in Fig 2d), is likely due to design imperfections in our cavities when introducing the thin Ti/Au layer. These design issues can be solved by more carefully compensating for SAW phase-velocity changes under the Ti/Au layer. In addition, we estimate the single-phonon coupling rate $g_{0}$ to be approximately 42 kHz for this device (Supplementary Information). Compared to the best $g_{0}$ achieved in previous work [46], this smaller $g_{0}$ is a result of larger cavity mode volume, which was used for easier characterization. The design principle described in this work can be immediately applied to the tightly focusing cavities studied in Ref. [46] to achieve $g_{0}$ larger than 1 MHz. Therefore, the integration of gated QDs and SAW resonators can be done without sacrificing the performance of individual components.
Discussion